# xTED: Cross-Domain Adaptation via Diffusion-Based Trajectory Editing

### v2

@@ 4.1 Model Architecture
Encoding and decoding design for decision making data.
States, actions, and rewards in MDP present fundamentally dissimilar notions, entailing distinct physical meanings and consistency properties. This naturally distinguishes the trajectory-based models from image-based models, which only considers spatial representations among visual patches (Peebles & Xie, 2023) and simply treats every pixel homogeneously in modeling designs (Ho et al., 2020). Thus, xTED encodes and decodes state, action, and reward sequences $\tau^{s},\tau^{a},\tau^{r}$ in trajectories $\tau$ separately. In every noise step $k$, each sequence is first compressed or enlarged into latent representation $h_{k}^{s},h_{k}^{a},h_{k}^{r}$ with appropriate feature dimensions using separate networks $f$, and then processed with a self-attention module S-Attn to capture global consistency along the sequence dimension (conditions are also embedded with $\mathbf{e}_{k}^{c}=f_{c}(k,\mathbf{c})$):
@@
The series of networks $f$ provides substantial flexibility to incorporate task-specific prior knowledge, facilitating the construction of a reasonable information bottleneck. The embedding spaces should not be overly compressed, as subsequent attention blocks require sufficiently rich information to uncover dependencies. Conversely, these embedding spaces must not be excessively enlarged due to concerns about computational complexity and memory usage. From an interactive perspective, the embedding dimensions should be balanced to prevent the model from neglecting lower-dimensional action and reward embeddings when processing high-dimensional visual observations. At the end, given the final representation $h_{k}^{\prime s},h_{k}^{\prime a},h_{k}^{\prime r}$, we use separate decoders to predict the corresponding noise $y_{k}$ at noise step $k$, which recover the original dimension space respectively:
$y_{k}^{s}=\texttt{D}_{\texttt{s}}(h_{k}^{\prime s})\in\mathbb{R}^{\mathcal{S}},\ y_{k}^{a}=\texttt{D}_{\texttt{a}}(h_{k}^{\prime s})\in\mathbb{R}^{\mathcal{A}},\ y_{k}^{r}=\texttt{D}_{\texttt{r}}(h_{k}^{\prime r})\in\mathbb{R}^{R}$ (7)
Crucially, this separation on encoding and decoding preserves the inherent differences between states, actions, and rewards, aiding the model in recognizing clear dependencies on the latent representations $h_{k}$. Compared with prior works, these help avoid exploiting spurious correlations within decision-making data and sidesteps modeling the overall trajectory distribution directly (Janner et al., 2022; He et al., 2023), thereby enhancing sample efficiency and maximizing model expressiveness.
Dependency structure modeling.
Although states, actions, and rewards possess fundamentally different physical properties, significant interactive dependencies exist among them. Specifically, unlike the spatial representation of images, sequential decision data entails intricate dynamics patterns that consistently drive state updates, action generation, and reward acquisition throughout trajectories, which play an indispensable role in effective trajectory modeling. To this end, the state and action embeddings $h_{k}^{s}$ and $h_{k}^{a}$ are cross-attended to capture mutual dependencies by exchanging key-value pairs in multi-headed attention, modeling the effects of policy and transition state dynamics separately:
$h_{k}^{\prime s}=\texttt{X-Attn}(h_{k}^{s},h_{k}^{a},\mathbf{e}_{k}^{c}),\quad h_{k}^{\prime a}=\texttt{X-Attn}(h_{k}^{a},h_{k}^{s},\mathbf{e}_{k}^{c})$ (8)
However, it is important to note that rewards are naturally dependent on state-action sequences, while this dependency does not work in reverse due to their underlying causal relationships. Enforcing in-context optimality guidance could conflict with the primary goal of addressing domain gaps within trajectories. Therefore, we query the reward embeddings $h_{k}^{r}$ using concatenated state-action embeddings, without imposing unreasonable reward dependencies on state-action sequence modeling:
@@
Introducing prior knowledge of the causal dependencies among states, actions, and rewards provides necessary regularizations for capturing trajectory distributions. This dependency structure modeling module alleviates the overabundance of information and relationships, enabling the model to accurately capture essential dependency priors and efficiently handle long-term sequential modeling. Additionally, the sequence-level dependency structure modeling naturally decouples the hypothesized temporal causal relationships adopted in many sequential modeling approaches (Chen et al., 2021; Ajay et al., 2023a), offering flexibility in whether to incorporate temporal causality during modeling.
Incorporating external conditions.
@@
In addition to the noise step $k$, our model architecture accommodates various conditioning information $\mathbf{c}=[c_{i}]_{i=0}^{N}$, guiding trajectories toward desirable regions. For example, we can steer the entire trajectory modeling toward high-rewarding regions by conditioning on normalized trajectory returns $\mathbf{c}=[R(\tau)]$ (Ajay et al., 2023a), with specific implementation details outlined in Appendix B.1. Moreover, our model demonstrates superior synergy with return-conditioned modeling compared to naive architectures that concatenate states, actions, and rewards along temporal or feature dimensions as input, as analyzed in Section 5.6. This advantage arises from our separate encoding-decoding manner and designs of dynamics dependency modeling, which effectively recognize the one-way dependencies between reward and state-action sequences.
Applicability to diverse decision-making task settings.
Our architecture is compatible with multi-modal observations, including proprioceptive, exteroceptive (visual/point clouds), and language inputs, as long as modality-specific encoders are employed. From the perspective of trajectory composition, it offers flexibility in accommodating non-standard trajectories, which may include terminal flags, class labels, or preference labels for auxiliary module training to aid policy optimization. If these task-specific elements exhibit similarities with rewards, which have specific dependencies on states and/or actions, the architecture can be easily extended with minimal additional design complexities by incorporating extra encoders, decoders, and attention modules, akin to our reward-relevant design philosophies. Furthermore, practitioners can also omit the reward-relevant modules to fit non-reward (imitation learning) settings, as demonstrated in our real robot experiments (Section 5.4).
4.2 Cross-Domain Trajectory Editing
Training Diffusion Model on Target Data.
@@ 4.2 Cross-Domain Trajectory Editing
Given a target domain dataset, we first train the diffusion model to capture its trajectory distribution using Eq. 5, where $\mathbf{x}_{k}$ represents the noised trajectory $\tau_{k}$, and additional conditions $\mathbf{c}$, if any, are concatenated with the noise step $k$ as inputs. Particularly, we keep the initial transition unchanged in each $H$-horizon trajectory during training, serving as an anchor for enhanced stability in long-sequence modeling.
Editing Source Trajectories.
@@
In the forward process, we obtain noised trajectories $\tau_{k}$ from $\mathcal{N}(\tau_{k};\tau_{0},\sigma^{2}_{k}\mathbf{I})$, which retains information from the original source trajectory $\tau_{0}$. This procedure is governed by a ratio parameter $\kappa=\frac{k}{K}\in[0,1]$, determining the level of noise addition. A small $\kappa$ preserves much of the source trajectory information, while a large $\kappa$ diminishes it. The ideal selection principle for $\kappa$ aims to distort only fine-grained information, such as transition dynamics in the source data, while retaining valuable large-scale information, such as skill primitives inherent in the trajectories. Additionally, this noised initialization theoretically yields favorable results by solving the reverse stochastic differential equation. Thus, we could then denoise $\tau_{k}$ using the diffusion prior, successfully aligning the edited trajectories $\hat{\tau}_{0}$ with the target domain. The initial transition remains unchanged after each forward and denoising step, consistent with the training process, which is crucial for enhancing editing performance, as discussed in Appendix C.6. Experiments in Appendix C.2 indicate that setting $\kappa=0.5$ is universally optimal for incorporating the edited source data into policy learning, ensuring sufficient exploitation of source data while shifting its domain properties toward the target domain. After the editing process, we can safely integrate edited source data with target data for downstream policy learning using any algorithms at task specifications.
@@ 5 Experiments
In this section, we present empirical evaluations of the proposed cross-domain trajectory editing framework, xTED. We begin with detailing our experimental setups and the baselines for comparison. We then assess xTED against baseline methods in simulation (5.3) and real-robot tasks (5.4), overcoming various observation, dynamics, and morphology gaps within the source data. xTED consistently demonstrates advantages across different domain gaps and tasks compared to solely training on original target data or directly augmenting target data with source data (5.7), as well as its stability on various data quantity (5.5). We conclude with ablation studies on key design elements of our model architecture and provide evidence and analyses that echo with insights from methodology (5.6).
@@ 5.1 Experimental Environment Setups
Simulation Experiments.
We conduct simulation-based experiments using the MuJoCo physics simulator (Todorov et al., 2012). Specifically, we construct two source domains on Walker2d-v2 (WK) and HalfCheetah-v2 (HC), where we intentionally introduce dynamics and morphology gaps relative to the original target domains by modifying the physics modeling configurations: (1) Gravity: applying 2×2$\times$2 × gravitational acceleration in the simulation dynamics; (2) Friction: using 0.25/0.5×0.25/0.5$\times$0.25 / 0.5 × friction coefficient to make the agent harder to maintain balance; (3) Thigh Size: using 2×2$\times$2 × thigh size to introduce morphology gaps on the embodiment. We record the average return of Soft Actor-Critic (Haarnoja et al., 2018) policies trained in these source domains and provide a straightforward quantification of the impact of domain gaps on source policy learning in Appendix B.3.
For data acquirement and implementation details: (i) target data: we randomly select 20k transitions from the corresponding datasets in the standard offline RL benchmark D4RL (Fu et al., 2020). Specifically, we only consider the Medium (M), Medium Replay (MR) and Medium Expert (ME) datasets, as we hardly use a random or expert policy for system control in common real-world scenarios. (ii) source data: we collect trajectories of 20k transitions from the source domain with a SAC policy trained in the same domain. Please see Appendix B.1 for more implementation details.
@@
Real-Robot Experiments.
We conduct real-world experiments in robotic environments where target data is collected by the WidowX robot and source data is collected by the Airbot, for 100 trajectories respectively. We build three manipulation tasks: (1) Pick up a red cup and place it on a silver pan (Cup); (2) Pick up a duck and place it on a green plate (Duck); (3) Move a pot from right to left (Pot). As shown in Fig. 3, there are huge domain discrepancies on robot embodiments and camera viewpoints: two 7-DoF robots possess distinctive module shapes/lengths, embodiment masses and joint types; source and target configurations present different viewpoints and field of views from base and wrist cameras on two robots. Our objective is to incorporate the edited data from the Airbot to enhance policy performance on the WidowX robot. These task environments are highly stochastic, featuring randomly initialized object positions and poses, as well as distractors with various shapes, colors, and locations. Please see Appendix B.2 for more implementation details.
@@ 5.2 Baselines
5.3 Comparative Evaluation in Simulation Experiments
In Table 1, we present the comparative results of xTED and other baselines on the simulation-based experiments. We integrate our trajectory editing process with a downstream policy learning module, implemented using one of the widely-used state-of-the-art (SOTA) offline RL algorithms, Implicit Q-Learning (IQL) (Kostrikov et al., 2022). Notably, our method achieves the best or on-par performance in almost all tasks (18 out of 18) compared to the baselines. Specifically, we find that policy learning directly augmented with original source data can have negative effects compared to solely learning from target data in 5 out of 18 tasks, while augmenting with edited source data consistently shows improvement in all of the tasks. In WK-MR and HC-MR, edited source data helps improve the performance by over 50% and 20% respectively. The overall 16.4% improvement also shows consistent stability on different domain gaps and quality of target data. This likely results from the fact that biased dynamics in source trajectories can detrimentally affect policy learning, despite the increased data coverage and diversified behavior patterns introduced by these trajectories. Nevertheless, xTED effectively adjusts the trajectories to realistic dynamics and provides meaningful augmentation, rather than merely expanding the data distribution. xTED also demonstrates the ability to handle vairous gaps simultaneously using a single fixed diffusion model in Appendix C.9.
@@ 5.3 Comparative Evaluation in Simulation Experiments
Target Data Source Dynamics Src Tgt Tgt+Src $\Delta$ Tgt+Src (Edited) $\Delta$ Halfcheetah Med Gravity 8.4$\pm$6.3 39.5$\pm$2.4 41.1$\pm$1.3 +4.1% 40.6$\pm$2.0 +2.8% Friction 14.4$\pm$3.6 39.5$\pm$2.4 27.9$\pm$7.2 -29.4% 41.2$\pm$1.6 +4.3% Thigh Size -0.1$\pm$1.0 39.5$\pm$2.4 40.3$\pm$2.2 +2.0% 40.7$\pm$2.4 +3.0% Med-R Gravity 8.4$\pm$6.3 26.2$\pm$3.5 29.5$\pm$2.9 +12.6% 31.3$\pm$2.9 +19.5% Friction 14.4$\pm$3.6 26.2$\pm$3.5 23.2$\pm$4.0 -11.5% 31.8$\pm$3.1 +21.4% Thigh Size -0.1$\pm$1.0 26.2$\pm$3.5 29.0$\pm$3.7 +10.7% 33.0$\pm$3.0 +26.0% Med-E Gravity 8.4$\pm$6.3 40.1$\pm$2.9 40.4$\pm$3.4 +0.7% 43.8$\pm$3.6 +9.2% Friction 14.4$\pm$3.6 40.1$\pm$2.9 27.1$\pm$5.4 -32.4% 43.2$\pm$3.0 +7.7% Thigh Size -0.1$\pm$1.0 40.1$\pm$2.9 37.4$\pm$5.5 -6.7% 43.0$\pm$3.0 +7.2% Walker2d Med Gravity 16.6$\pm$6.2 45.3$\pm$15.9 60.4$\pm$9.5 +33.3% 58.2$\pm$11.7 +28.5% Friction 8.2$\pm$1.4 45.3$\pm$15.9 47.0$\pm$12.8 +3.8% 54.5$\pm$13.7 +20.3% Thigh Size 7.6$\pm$3.5 45.3$\pm$15.9 49.6$\pm$12.2 +9.5% 58.9$\pm$11.7 +30.0% Med-R Gravity 16.6$\pm$6.2 16.6$\pm$5.9 19.5$\pm$10.7 +17.5% 23.3$\pm$9.1 +40.3% Friction 8.2$\pm$1.4 16.6$\pm$5.9 17.4$\pm$6.2 +4.8% 25.9$\pm$9.1 +56.0% Thigh Size 7.6$\pm$3.5 16.6$\pm$5.9 18.0$\pm$6.7 +8.4% 25.9$\pm$9.1 +56.0% Med-E Gravity 16.6$\pm$6.2 71.0$\pm$21.0 67.4$\pm$11.1 -16.9% 82.9$\pm$18.1 +16.8% Friction 8.2$\pm$1.4 71.0$\pm$21.0 75.1$\pm$18.3 +5.8% 74.0$\pm$24.5 +4.2% Thigh Size 7.6$\pm$3.5 71.0$\pm$21.0 77.0$\pm$19.3 +8.5% 81.0$\pm$21.4 +14.1% Total 165.3 716.1 727.3 +1.6% 833.2 +16.4%
5.4 Comparative Evaluation in Real-Robot Experiments
The average success rates for Cup, Duck, and Pot tasks are shown in Fig. 3. Each result is estimated over 10 episodes per task with random environmental initializations and averaged over 3 seeds. Target + Edited Source overwhelmingly outperforms the baselines in all the tasks, regardless of the presence of distractors. Notably, in the Cup task, combining learning from edited source data increases the success rate from 43% to 97% and 30% to 73.3%, in environments without and with distractors respectively. In contrast, due to the huge domain gaps between the source and target domains, directly adding original source data yields no performance gain and sometimes even causes degradation against solely training on target data, e.g. 40% to 20% in Duck task and 23% to 0% in Pot task. When evaluated with surrounding distractors, directly incorporating source data into policy learning sometimes leads to a more severe performance drop than without distractors, such as from 43% to 13% and from 40% to 0% in the Duck and Pot tasks, respectively. Furthermore, our method consistently outperforms the Target with S4RL across all tasks. This result highlights that data augmentation methods cannot reliably enhance policy performance when compared to using only the original target data. It underscores the superior effectiveness and necessity of the cross-domain trajectory editing paradigm, which appropriately leverages biased source data. All numeric results in Fig. 3 can be found in Appendix C.1 Table 7.
5.5 Investigations on Sample Efficiency
Figure 4: Different amounts of data for diffusion and policy training (Target: WK-MR, Source: Gravity).
xTED demonstrates robust performance across varying amounts of data. We first ablate on data quantity for diffusion training (5k, 10k, 20k transitions from WK-MR) and keep the same data quantity for policy learning. From Table 2, we find that the policies learned from target and edited source data still perform well despite of very limited data for training diffusion model, demonstrating high sample efficiency of diffusion models in our data editing paradigm.
We then ablate on data quantity for both diffusion training and policy learning: randomly sample transitions of 50k and 100k from the WK-MR dataset for diffusion training, combining each with an equivalent amount of source data for policy learning. Fig. 4 exhibits the policy performance on target, target+source, and target+edited source datasets, utilizing different training algorithms, IQL and TD3+BC (Fujimoto & Gu, 2021). It is evident that with small datasets (20k transitions, Table 1), our method significantly enhances the quality of the source domain dataset, thereby improving training performance. As the dataset size increases (50k and 100k transitions), our method continues to enhance source data quality, though the improvement over baselines might sometimes become less pronounced since the target data itself can sufficiently support policy learning without source data. xTED also prevails on larger amounts of source data involved in policy learning in Appendix C.4.
Table 2: Reduced data usage for diffusion training hardly has negative impact on policy learning.
Domain Gap Gravity Friction Thigh Size
Target Data 5k 10k 20k 5k 10k 20k 5k 10k 20k
Avg. Score 26.7±9.8 21.5±8.5 23.3±9.1 29.0±8.3 25.0±8.2 25.9±9.1 31.4±10.0 29.8±9.6 25.9±9.1
5.6 Ablation Findings on Model Architecture

### v3

@@ 4.1 Model Architecture
Encoding and decoding design for trajectory data. States, actions, and rewards in MDP present fundamentally distinct notions and physical meanings. This naturally distinguishes the trajectory-based models from image-based models, which only consider spatial representations among visual patches (Peebles & Xie, 2023) and simply treat every pixel homogeneously in modeling designs (Ho et al., 2020). Thus, xTED encodes and decodes state, action, and reward sequences $\tau^{s},\tau^{a},\tau^{r}$ in trajectories $\tau$ separately, which offers great flexibility to incorporate task-specific prior knowledge for each element. In every noise step $k$, each sequence is first encoded into latent representation $h_{k}^{s},h_{k}^{a},h_{k}^{r}$ with using separate sub-networks $f$, and then processed with a self-attention block S-Attn for temporal modeling along the sequence dimension (conditions are also embedded with $\mathbf{e}_{k}^{c}=f_{c}(k,\mathbf{c})$, short for $\mathbf{e}$ in Fig. 2):
@@
At the end, given the final representation $h_{k}^{\prime s},h_{k}^{\prime a},h_{k}^{\prime r}$, we use separate decoders to predict the corresponding noise $y_{k}$ at noise step $k$, which recover the original data dimension:
$y_{k}^{s}=\texttt{D}_{\texttt{s}}(h_{k}^{\prime s})\in\mathbb{R}^{H\times|\mathcal{S}|},\ y_{k}^{a}=\texttt{D}_{\texttt{a}}(h_{k}^{\prime s})\in\mathbb{R}^{H\times|\mathcal{A}|},$
$y_{k}^{r}=\texttt{D}_{\texttt{r}}(h_{k}^{\prime r})\in\mathbb{R}^{H\times 1}$
Crucially, this separation on encoding and decoding preserves the inherent differences between states, actions, and rewards, aiding the model in recognizing clear temporal dependencies on the latent representations $h_{k}$. Compared with prior works that concatenate and process the state, action, and reward as an image-like big matrix (Janner et al., 2022; He et al., 2023), our design avoids exploiting spurious correlations within decision-making data and sidesteps modeling the joint trajectory distribution directly, thereby enhancing sample efficiency and maximizing model expressiveness.
Dependency structure modeling. Despite the distinction across states, actions, and rewards, significant internal dependencies exist. Specifically, unlike the 2D representation of images, the transition dynamics and reward function pose strong coupling and intricate internal dependency within sequential decision-making data. To this end, the state and action embeddings $h_{k}^{s}$ and $h_{k}^{a}$ are cross-attended to capture mutual dependencies between states and actions by exchanging key-value pairs in multi-headed attention:
$h_{k}^{\prime s}=\texttt{X-Attn}(h_{k}^{s},h_{k}^{a},\mathbf{e}_{k}^{c}),\ h_{k}^{\prime a}=\texttt{X-Attn}(h_{k}^{a},h_{k}^{s},\mathbf{e}_{k}^{c})$ (7)
Also, it is important to note that rewards are naturally dependent on state-action pairs, while this dependency does not work in the reverse direction due to their underlying causal relationships. Therefore, we query the reward embeddings $h_{k}^{r}$ using concatenated state-action embeddings, without imposing unreasonable reward dependencies on state-action sequence modeling:
@@
Introducing prior knowledge of the causal dependencies among states, actions, and rewards facilitates accurately capturing essential dependencies and temporal consistency within trajectory data, therefore improving sample efficiency and providing necessary regularizations when modeling complex trajectory distributions.
@@
Incorporating external conditions.  In addition to noise step $k$, our model architecture accommodates various conditioning information $\mathbf{c}=[c_{i}]_{i=0}^{N}$, guiding trajectories toward desirable patterns. For example, we can steer the entire trajectory modeling toward high-rewarding regions by conditioning on normalized trajectory returns $\mathbf{c}=[R(\tau)]$ (Ajay et al., 2023a), with implementation details outlined in Appendix B.1. Moreover, our model demonstrates superior synergy with return-conditioned modeling compared to naïve architectures that concatenate states, actions, and rewards along temporal or feature dimensions as input, as analyzed in Appendix C.6. This advantage arises from our separate encoding-decoding design and dependency structure modeling, which effectively recognize the causal dependencies between reward and state-action sequences.
4.2 Cross-Domain Trajectory Editing with Diffusion
@@
Training diffusion model on target data. Given a target domain dataset, we first train our proposed diffusion model architecture in Fig. 2 to capture its trajectory distribution using diffusion loss Eq. (5), where $\mathbf{x}_{k}$ represents the noised trajectory $\tau_{k}$. Additional conditions $\mathbf{c}$, if any, are concatenated with the noise step $k$ as inputs. Particularly, we keep the initial transition unchanged in each $H$-horizon trajectory during training, serving as an anchor for enhanced stability in long-sequence modeling.
Editing source trajectories. The editing process contains two stages: the forward process, which perturbs source data with noise, and the reverse process, which aligns noised source data with target domain properties.
@@
In the forward process, we obtain noised source trajectories $\tilde{\tau}_{k}$ from $\mathcal{N}(\tilde{\tau}_{k};\tilde{\tau}_{0},\sigma^{2}_{k}\mathbf{I})$, which retains some information from the original source trajectories $\tilde{\tau}_{0}$. This procedure is governed by a ratio parameter $\kappa=\frac{k}{K}\in[0,1]$, determining the level of noise addition. A small $\kappa$ preserves much of the source trajectory information, while a large $\kappa$ diminishes it. The ideal selection principle for $\kappa$ is to blur only fine-grained data details, such as transition dynamics in the source data, while retaining valuable mesoscopic information, such as skill primitives in the source trajectories. This noised initialization can yield favorable results by solving the reverse stochastic differential equation, which has achieved past success in diffusion-based image editing (Meng et al., 2022; Kawar et al., 2023; Ceylan et al., 2023).
In the reverse process, we denoise the noised $\tilde{\tau}_{k}$ using the learned target diffusion prior, aligning the edited trajectories $\hat{\tau}_{0}$ with the target domain. The initial state-action-reward triple is kept unchanged during the editing process, consistent with the training process, which is crucial for enhancing editing performance as discussed in Appendix C.6. Our empirical tests (see Appendix C.2) indicate that setting $\kappa=0.5$ is practically optimal for incorporating the edited source data into policy learning, which ensures sufficient exploitation of source data while shifting its domain properties toward the target domain.
4.3 Applicability to Diverse Downstream Task Settings
The xTed framework can serve as an extra data pre-processing step with great flexibility. The edited source data can be integrated with target data for downstream policy learning using any algorithms (IL or RL, single- or cross-domain). Our proposed diffusion architecture is also compatible with multi-modal observations, including proprioceptive, exteroceptive (visual/point clouds), and language inputs, as long as modality-specific encoders are employed. If these task-specific elements exhibit similarities with rewards, which have specific dependencies on states and/or actions, the architecture can be easily extended with minimal additional design complexities by incorporating extra encoders, decoders, and attention blocks, akin to our reward dependency design. Furthermore, practitioners can also omit the reward-relevant modules to fit non-reward (imitation learning) settings, as demonstrated in our real robot experiments (Section 5.3).
@@ 5 Experiments
In this section, we present empirical evaluations of xTED. We begin with detailing our experimental setups and the baselines for comparison. We combine xTED with various single-domain and cross-domain IL/RL methods, and compare against baseline methods in simulation and real-robot tasks (Section 5.3), overcoming various observation, dynamics, and morphology gaps within the source data. xTED consistently demonstrates superior performance across different domain gaps and tasks compared to solely training on original target data or directly augmenting target data with source data (Section 5.4). We conclude with ablation studies on key design elements of our model architecture and provide evidence and analyses that echo with insights from methodology (Section 5.5).
@@ 5.1 Experimental Environment Setups
Real-robot experiments. We conduct real-world experiments in robotic environments where target data is collected by the WidowX robot and source data is collected by the Airbot, for 100 trajectories respectively. We build three manipulation tasks: (1) Pick up a red cup and place it on a silver pan (Cup); (2) Pick up a duck and place it on a green plate (Duck); (3) Move a pot from right to left (Pot). As shown in Fig. 3, there are huge domain discrepancies on robot embodiments and camera viewpoints: two 7-DoF robots possess distinctive module shapes/lengths, embodiment masses and joint types; source and target configurations present different viewpoints and field of views from base and wrist cameras on two robots. Our objective is to incorporate the edited data from the Airbot to enhance policy performance on the WidowX robot. These task environments are highly stochastic, featuring randomly initialized object positions and poses, as well as distractors with various shapes, colors, and locations. Please see Appendix B.2 for more implementation details.
Simulation experiments. We conduct simulation-based experiments using the MuJoCo physics simulator (Todorov et al., 2012). Specifically, we construct two source domains on Walker2d-v2 (WK) and HalfCheetah-v2 (HC), where we intentionally introduce dynamics and morphology gaps relative to the original target domains by modifying the physics modeling configurations: (1) Gravity: applying 2×2$\times$2 × gravitational acceleration in the simulation dynamics; (2) Friction: using 0.25/0.5×0.25/0.5$\times$0.25 / 0.5 × friction coefficient to make the agent harder to maintain balance; (3) Thigh Size: using 2×2$\times$2 × thigh size to introduce morphology gaps on the embodiment. We record the average return of Soft Actor-Critic (Haarnoja et al., 2018) policies trained in these source domains and provide a straightforward quantification of the impact of domain gaps on source policy learning in Appendix B.3.
For data acquirement and implementation details: (i) target data: we randomly select 20k transitions from the corresponding datasets in the standard offline RL benchmark D4RL (Fu et al., 2020). Specifically, we only consider the Medium (M), Medium Replay (MR) and Medium Expert (ME) datasets, as we hardly use a random or expert policy for system control in common real-world scenarios. (ii) source data: we collect trajectories of 20k transitions from the source domain with a SAC policy trained in the same domain. See Appendix B.1 for more implementation details.
@@ 5.2 Baselines
5.3 xTED with Various Policy Learning Methods
xTED+single-domain IL. For real-robot experiments, we edit source data with xTED and then learn policies with behavior cloning (BC). The average success rates for Cup, Duck, and Pot tasks are shown in Fig. 3. Each result is estimated over 10 episodes per task with random environmental initializations and averaged over 3 seeds. Target + Edited Source overwhelmingly outperforms the baselines in all the tasks, regardless of the presence of distractors. Notably, in the Cup task, combining learning from edited source data increases the success rate from 43% to 97% and 30% to 73.3%, in environments without and with distractors respectively. In contrast, due to the huge domain gaps between the source and target domains, directly adding original source data yields no performance gain and sometimes even causes degradation against solely training on target data, e.g. 40% to 20% in Duck task and 23% to 0% in Pot task. When evaluated with surrounding distractors, directly incorporating source data into policy learning sometimes leads to a more severe performance drop than without distractors, such as from 43% to 13% and from 40% to 0% in the Duck and Pot tasks, respectively. Furthermore, our method consistently outperforms the Target with S4RL across all tasks. This result highlights that data augmentation methods cannot reliably enhance policy performance when compared to using only the original target data. It underscores the superior effectiveness and necessity of the cross-domain trajectory editing paradigm, which appropriately leverages biased source data. All numeric results in Fig. 3 can be found in Appendix C.1 Table 6.
xTED+single-domain RL. In Table 1, we present the comparative results of xTED and other baselines on the simulation-based experiments. We integrate our trajectory editing process with one of the widely-used state-of-the-art (SOTA) offline RL algorithms, Implicit Q-Learning (IQL) (Kostrikov et al., 2022). Notably, our method achieves the best or on-par performance in almost all tasks (18 out of 18) compared to the baselines. Specifically, we find that policy learning directly augmented with original source data can have negative effects compared to solely learning from target data in 5 out of 18 tasks, while augmenting with edited source data consistently shows improvement in all of the tasks. In WK-MR and HC-MR, edited source data helps improve the performance by over 50% and 20% respectively. The overall 16.4% improvement also shows consistent stability on different domain gaps and quality of target data. This likely results from the fact that biased dynamics in source trajectories can detrimentally affect policy learning, despite the increased data coverage and diversified behavior patterns introduced by these trajectories. Nevertheless, xTED effectively adjusts the trajectories to realistic dynamics and provides meaningful augmentation, rather than merely expanding the data distribution. xTED also demonstrates the ability to handle various gaps simultaneously using a single fixed diffusion model in Appendix C.9.
xTED+cross-domain methods. xTED can be integrated with any cross-domain IL/RL method to further enhance performance, as xTED helps reduce domain gaps at data level and ease the burden of downstream cross-domain policy learning. To evaluate this, we combine xTED with DARA (Liu et al., 2022) that corrects rewards in source data with domain discriminators. Specifically, we implement DARA on offline target and source data edited by xTED. The results in Table 2 demonstrate that xTED consistently improves DARA’s performance since xTED helps increase the overlap of source and target distribution which could facilitate domain discriminator training in DARA. Similar to DARA, many cross-domain policy adaptation methods rely on careful balance of source and target data or/and sufficient data availability to ensure effective domain correspondence/discriminator learning. Applying xTED for data pre-processing could reduce domain heterogeneity in target and source data, relaxing these requirements and enabling more data-efficient and robust cross-domain adaptation.
@@
Target Data Domain Gap Src Tgt Tgt+Src $\Delta$ Tgt+Src (Edited) $\Delta$ Halfcheetah Med Gravity 8.4$\pm$6.3 39.5$\pm$2.4 41.1$\pm$1.3 +4.1% 40.6$\pm$2.0 +2.8% Friction 14.4$\pm$3.6 39.5$\pm$2.4 27.9$\pm$7.2 -29.4% 41.2$\pm$1.6 +4.3% Thigh Size -0.1$\pm$1.0 39.5$\pm$2.4 40.3$\pm$2.2 +2.0% 40.7$\pm$2.4 +3.0% Med-R Gravity 8.4$\pm$6.3 26.2$\pm$3.5 29.5$\pm$2.9 +12.6% 31.3$\pm$2.9 +19.5% Friction 14.4$\pm$3.6 26.2$\pm$3.5 23.2$\pm$4.0 -11.5% 31.8$\pm$3.1 +21.4% Thigh Size -0.1$\pm$1.0 26.2$\pm$3.5 29.0$\pm$3.7 +10.7% 33.0$\pm$3.0 +26.0% Med-E Gravity 8.4$\pm$6.3 40.1$\pm$2.9 40.4$\pm$3.4 +0.7% 43.8$\pm$3.6 +9.2% Friction 14.4$\pm$3.6 40.1$\pm$2.9 27.1$\pm$5.4 -32.4% 43.2$\pm$3.0 +7.7% Thigh Size -0.1$\pm$1.0 40.1$\pm$2.9 37.4$\pm$5.5 -6.7% 43.0$\pm$3.0 +7.2% Walker2d Med Gravity 16.6$\pm$6.2 45.3$\pm$15.9 60.4$\pm$9.5 +33.3% 58.2$\pm$11.7 +28.5% Friction 8.2$\pm$1.4 45.3$\pm$15.9 47.0$\pm$12.8 +3.8% 54.5$\pm$13.7 +20.3% Thigh Size 7.6$\pm$3.5 45.3$\pm$15.9 49.6$\pm$12.2 +9.5% 58.9$\pm$11.7 +30.0% Med-R Gravity 16.6$\pm$6.2 16.6$\pm$5.9 19.5$\pm$10.7 +17.5% 23.3$\pm$9.1 +40.3% Friction 8.2$\pm$1.4 16.6$\pm$5.9 17.4$\pm$6.2 +4.8% 25.9$\pm$9.1 +56.0% Thigh Size 7.6$\pm$3.5 16.6$\pm$5.9 18.0$\pm$6.7 +8.4% 25.9$\pm$9.1 +56.0% Med-E Gravity 16.6$\pm$6.2 71.0$\pm$21.0 67.4$\pm$11.1 -16.9% 82.9$\pm$18.1 +16.8% Friction 8.2$\pm$1.4 71.0$\pm$21.0 75.1$\pm$18.3 +5.8% 74.0$\pm$24.5 +4.2% Thigh Size 7.6$\pm$3.5 71.0$\pm$21.0 77.0$\pm$19.3 +8.5% 81.0$\pm$21.4 +14.1% Total 165.3 716.1 727.3 +1.6% 833.2 +16.4%
Table 2: Validation on xTED combined with cross-domain policy learning method DARA. Average normalized scores for MuJoCo tasks on 20k target/source data (5 random seeds).